a pizza is cut into 8 equal slices.if the radius of the pie is 12 inches,what is the area of 2 of the slices
Found 2 solutions by longjonsilver, cleomenius:
Answer by longjonsilver(2297)

i have answered a duplicate version of this

Answer by cleomenius(959)

Area = x/360 ( pi * r^2)
2/8 = 1/4
1/4( 3.14 144) = ~113
A ~ 113 in ^2